### v1

@@ Abstract
Various tasks in scientific computing can be modeled as an optimization problem on the indefinite Stiefel manifold. We address this using the Riemannian approach, which basically consists of equipping the feasible set with a Riemannian metric, preparing geometric tools such as orthogonal projections, formulae for Riemannian gradient, retraction and then extending an unconstrained optimization algorithm on the Euclidean space to the established manifold. The choice for the metric undoubtedly has a great influence on the method. In the previous work [D.V. Tiep and N.T. Son, A Riemannian gradient descent method for optimization on the indefinite Stiefel manifold, arXiv:2410.22068v2[math.OC]], a tractable metric, which is indeed a family of Riemannian metrics defined by a symmetric positive-definite matrix depending on the contact point, has been used. In general, it requires solving a Lyapunov matrix equation every time when the gradient of the cost function is needed, which might significantly contribute to the computational cost. To address this issue, we propose a new Riemannian metric for the indefinite Stiefel manifold. Furthermore, we construct the associated geometric structure, including a so-called quasi-geodesic and propose a retraction based on this curve. We then numerically verify the performance of the Riemannian gradient descent method associated with the new geometry and compare it with the previous work.
@@ 1 Introduction
Optimization of smooth functions of matrix variable $X$ subject to the constraint of the form $\mathcal{M}=\{X\ :\ X^{T}AX=J\}$, where $A$ is nonsingular and symmetric and $J$ is symmetric, involutory, i.e., satisfying $J^{2}=I$, is an established line of researches due to its wide range of applications. Various problems such as dimensionality reduction of data via principal component analysis, canonical correlation analysis, finding locality preserving projection, and computation of the generalized eigenvalues of a positive-definite matrix pencil can be formulated as an optimization problem with such a constraint, see, e.g., HardSST04 ; HeN03 ; Hote92 ; KovaV95 ; SameW82 ; TiepSon2025 ; YgerBGR12 .
Additional conditions can be imposed on the matrices $A$ and $J$ upon the applications considered making the feasible set a differentiable manifold. When $A=I_{n}$ and $J=I_{k},k\leq n$, the $n\times n$ and $k\times k$ identity matrices, respectively, $\mathcal{M}$ is termed the (orthogonal) Stiefel manifold EdelAS98
@@
see, e.g., ShusA2023 . Although $\mathrm{St}(k,n)$ is a special case of $\mathrm{St}_{A}(k,n)$, one can see that $\mathrm{St}_{A}(k,n)$ can be turned into a Stiefel manifold by the mapping $X\mapsto R^{-1}Y$ where $A=R^{T}R$ is the Cholesky factorization of $A$. A more general one in which $A$ is nonsingular, symmetric, and $J$ is symmetric, involutory, which covers the orthogonal and generalized Stiefel manifolds as special cases, is the indefinite Stiefel manifold
@@
This manifold was recently introduced in TiepSon2025 . Exploiting the rich geometry of the feasible set, a popular approach, known as Riemannian optimization, is to extend well-known methods for unconstrained optimization in the Euclidean space to the case of these Riemannian manifolds once necessary geometric tools such as a Riemannian metric, orthogonal projections, a retraction are constructed. One is referred to classical works AdleDMMS02 ; EdelAS98 , monographs AbsiMS08 ; Boumal23 ; Sato2021 ; Udri1994 , and articles manopt ; MishS2016 ; SatoA19 ; ShusA2023 ; TiepSon2025 ; WangDPY2024 ; YgerBGR12 , to name a few among many others dedicated to optimization methods on the orthogonal, generalized, and indefinite Stiefel manifolds.
Equipping a Riemannian metric for a differentiable manifold is an essential step in Riemannian optimization. Recall that a Riemannian metric is an inner product defined on each tangent space of the manifold depending smoothly on the contact point. It not only helps quantify objects such as the length of vectors, the measure of angles but also has a great effect on the behavior of numerical methods. Specifically, it is a means of defining the Riemannian gradient of a cost function, a crucial material in most of optimization methods.
There are different ways to equip a manifold with a Riemannian metric depending on how the manifold is viewed. For some left/right-invariant Riemannian metrics, pseudo-Riemannian metrics equipped for the Stiefel manifold as a quotient manifold of two orthogonal groups, one is referred to (EdelAS98, , Subsect. 2.3), JurdML2020 ; NishA2005 ; cf. BendZ21 for the symplectic Stiefel manifold. Here, we follow (TiepSon2025, , Prop. 2.2) in which $\mathrm{iSt}_{A,J}(k,n)$ is shown to be an embedded submanifold of $\mathbb{R}^{n\times k}$} of dimension $nk-\frac{1}{2}k(k+1)$. Then, a Riemannian metric is defined via an $n\times n$ spd $X$-smoothly depending matrix $\mathbf{M}_{X}$ in which the inner product of $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}_{A,J}(k,n)$ is determined as
@@
This is not the most general way of defining a metric on a submanifold of $n\times k$ structured matrices but it covers most cases in practice, e.g., the standard Euclidean metric and the canonical metric EdelAS98 . Using this formulation, the Riemannian gradient of smooth cost functions via orthogonally projecting onto the tangent space requires solving a Lyapunov matrix equation (TiepSon2025, , Prop. 3.3), which can be computationally expensive when $k$ is not small. It is surprising that in the case of Stiefel manifold, the Riemannian gradient with respect to the Euclidean and canonical metrics have closed-form expressions, see (EdelAS98, , Sect. 2) and (AbsiMS08, , Subsect. 3.6). Moreover, in the generalized Stiefel manifold $\mathrm{St}_{A}(k,n)$ when $A$ is chosen to be the representing matrix $\mathbf{M}_{X}$, a closed-form Riemannian gradient is available (WangDPY2024, , Sect. 4). Meanwhile, for the symplectic Stiefel manifold, a relatively similar manifold to $\mathrm{St}_{A}(k,n),$ computing the Riemannian gradient needs solving a Lyapunov matrix equation under the Euclidean metric and a general tractable metric and does not when the canonical-like metric is used GSAS21a ; GSAS21 ; GSS2024_2 . On the one hand, these facts raise a hope that one can design a Riemannian metric, most probably the canonical(-like) metric as in EdelAS98 ; GSAS21 , for the indefinite Stiefel manifold such that solving the Lyapunov matrix equation is not compulsory in order to compute the Riemannian gradient. Note, however, that the motivation for using the canonical(-like) metric in EdelAS98 ; GSAS21 is to balance the contribution of the two components of tangent vectors in the metric formulation instead of obviating the necessity of solving the Lyapunov matrix equation as that for the metric proposed in this work. On the other hand, they inspire several questions: (1) How are we on track for not having to solve the Lyapunov matrix equations to compute the Riemannian gradient, (2) How to explain the different situations for orthogonal, generalized, and symplectic Stiefel manifolds mentioned above, and (3) How does an optimization method such as Riemannian gradient descent behave using the geometric quantities under the proposed metric?
@@ 2 Preliminaries
    where $W=JX^{T}AZ,K=(X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}Z,$ and $A^{-1}X_{\perp}K$ is independent of $X_{\perp}.$
@@
In (TiepSon2025, , Sect. 3), a so-called tractable metric $\mathrm{g}_{\mathbf{M}_{X}}$ defined via an $X$-smoothly dependent matrix $\mathbf{M}_{X}\in\mathrm{SPD}(n)$ was proposed. Namely, for $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}(k,n),$ $\langle Z_{1},Z_{2}\rangle_{\mathrm{g}_{\mathbf{M}_{X}}}\coloneq\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right)$. With this metric, for any $X\in\mathrm{iSt}(k,n)$, the ambient space $\mathbb{R}^{n\times k}$} is decomposed as the direct sum of the tangent space and the normal space,
@@ 3 A generalized canonical metric on $\mathrm{iSt}(k,n)$
For a greater detail, let us convert this matrix equation into a linear equation using vectorization operator and Kronecker product HornJ91 . Namely, for a matrix $M=(m_{ij})\in{\mathbb{R}^{n\times k}}$, $\mathrm{vec}(M)\in\mathbb{R}^{nk}$ is generated by stacking the columns of $M$ to make a column vector and the Kronecker product of $M$ and $N\in\mathbb{R}^{p\times q}$, denoted by $M\otimes N\in\mathbb{R}^{np\times kq}$, is the matrix consisting of $nk$ block matrices $m_{ij}N$ for $i=1,\ldots,n,j=1,\ldots,k$. In view of (HornJ91, , Lem. 4.3.1), (11) is equivalent to
@@
We say that a square system of linear equations is solving-free if its coefficient is a scalar matrix, i.e., the product of a (nonzero) scalar and the identity matrix. Thus, its solution is merely the product of the reciprocal of the scalar and the right-hand side. Accordingly, a sufficient condition for obviating the need of solving the Lyapunov matrix equation (8) is that $I_{k}\otimes(J\Gamma_{1}J)+(J\Gamma_{1}J)\otimes I_{k}$ has the form $2\rho I_{k^{2}}$, where $\rho$ is a nonzero constant parameter. It follows that $J\Gamma_{1}J=\rho I_{k}$ which, together with the condition $J^{2}=I_{k}$, yields $\Gamma_{1}=\rho I_{k}$. For the symmetric positive-definiteness of $\mathbf{M}_{X}^{-1}$, we set $\rho>0$. One can verify that in this case, the solution to (11) is $U_{X,Y}=\frac{1}{\rho}\mathrm{sym}(X^{T}AY).$ Furthermore, we set $\Gamma_{2}=0$ for the ease of argument. Consequently, $\mathbf{M}_{X}^{-1}$ takes the form
@@
Moreover, taking (3) into account, we also have
@@
Obviously, $g_{\rho,X_{\perp}}$ is a non-degenerate bilinear form defined on $T_{X}\mathrm{iSt}(k,n)$ and dependent upon both $\rho$ and $X_{\perp}$. To make this a Riemannian metric, except for the special parameter $\rho$, it is essential to require that $\mathbf{M}_{X}$ depends only on $X\in\mathrm{iSt}(k,n)$ and this dependence is smooth. To this end, we will impose additional conditions on $X_{\perp}$ in terms of $\Gamma_{3}$. If we choose
@@
  • One can see that when $\rho=2$ and $\Gamma_{3}=I_{n-k}$, the formulation (16) reduces to the canonical metric given in (EdelAS98, , Eq. (2.39)). Hence, the name generalized canonical metric. Moreover, we use this value for $\rho$ in numerical tests unless stated otherwise.
@@ 3.1 Geometry under the generalized canonical metric
For any $X\in\mathrm{iSt}(k,n),$ the matrix $\mathbf{M}_{X}$ defined by (13) is symmetric positive-definite, so we can extend the inner product $g_{\rho,X_{\perp}}$ on $T_{X}\mathrm{iSt}(k,n)$ to an inner product on $\mathbb{R}^{n\times k}$} which is given by $g_{\rho,X_{\perp}}(Z_{1},Z_{2})=\mathrm{tr}(Z_{1}^{T}\mathbf{M}_{X}Z_{2}),$ for all $Z_{1},Z_{2}\in{\mathbb{R}^{n\times k}}.$ The normal space to $\mathrm{iSt}(k,n)$ at $X$ with respect to $g_{\rho,X_{\perp}}$ is defined by
@@ 3.2 Riemannian gradient
Proof In view of the definition of the Riemannian gradient and (14), there holds
@@
for any $Z\in{\mathbb{R}^{n\times k}}.$ Therefore, from (12),
@@
In the case that $\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}=\Gamma_{3}$, taking (5) into account, it follows that
@@
If matrix $A$ is dense and has no special structure, the computational cost for $A^{-1}(\cdot)$ in (23) is generally comparable to $n^{3}$. Therefore in this case, using (24) is more advantageous. One can see that the computation of (24) requires $2n^{2}k+10nk^{2}-nk-3k^{2}$ flops. Obviously, when $k$ is approaching $n$, this cost is cheaper than that caused by (9). This will be seen in the numerical examples presented in Section 6.
@@
This explanation is also applicable to the geometry of the symplectic Stiefel manifold under the canonical-like metric GSAS21 as a special case of the tractable metric GSS2024_2 .

### v2

@@ Abstract
Various tasks in scientific computing can be modeled as an optimization problem on the indefinite Stiefel manifold. We address this using the Riemannian approach, which basically consists of equipping the feasible set with a Riemannian metric, preparing geometric tools such as orthogonal projections, formulae for Riemannian gradient, retraction and then extending an unconstrained optimization algorithm on the Euclidean space to the established manifold. The choice for the metric undoubtedly has a great influence on the method. In the previous work [D.V. Tiep and N.T. Son, A Riemannian gradient descent method for optimization on the indefinite Stiefel manifold, arXiv:2410.22068v2[math.OC]], a tractable metric, which is indeed a family of Riemannian metrics defined by a symmetric positive-definite matrix depending on the contact point, has been used. In general, it requires solving a Lyapunov matrix equation every time the gradient of the cost function is needed, which might significantly contribute to the computational cost. To address this issue, we propose a new Riemannian metric for the indefinite Stiefel manifold. Furthermore, we construct the associated geometric structure, including a so-called quasi-geodesic and propose a retraction based on this curve. We then numerically verify the performance of the Riemannian gradient descent method associated with the new geometry and compare it with the previous work.
@@ 1 Introduction
Optimization of smooth functions of matrix variable $X$ subject to the constraint of the form $\mathcal{M}=\{X\ :\ X^{T}AX=J\}$, where $A$ is nonsingular and symmetric, and $J$ is symmetric, involutory, i.e., satisfying $J^{2}=I$, is an established line of research due to its wide range of applications. Various problems such as dimensionality reduction of data via principal component analysis, canonical correlation analysis, finding locality preserving projection, and computation of the generalized eigenvalues of a positive-definite matrix pencil can be formulated as an optimization problem with such a constraint; see, e.g., HardSST04 ; HeN03 ; Hote92 ; KovaV95 ; SameW82 ; TiepSon2025 ; YgerBGR12 .
Additional conditions can be imposed on the matrices $A$ and $J$ upon the applications considered making the feasible set a differentiable manifold. When $A=I_{n}$ and $J=I_{k},k\leq n$, the $n\times n$ and $k\times k$ identity matrices, respectively, $\mathcal{M}$ is termed the orthogonal Stiefel manifold EdelAS98
@@
see, e.g., ShusA2023 . Although $\mathrm{St}(k,n)$ is a special case of $\mathrm{St}_{A}(k,n)$, one can see that $\mathrm{St}_{A}(k,n)$ can be turned into an orthogonal Stiefel manifold by the mapping $X\mapsto R^{-1}Y$ where $A=R^{T}R$ is the Cholesky factorization of $A$. A more general one in which $A$ is nonsingular, symmetric, and $J$ is symmetric, involutory, which covers the orthogonal and generalized Stiefel manifolds as special cases, is the indefinite Stiefel manifold
@@
This manifold was recently introduced in TiepSon2025 . Exploiting the rich geometry of the feasible set, a popular approach, known as Riemannian optimization, is to extend well-known methods for unconstrained optimization in the Euclidean space to the case of these Riemannian manifolds once necessary geometric tools such as a Riemannian metric, an orthogonal projection, and a retraction are constructed. One is referred to classical works AdleDMMS02 ; EdelAS98 , monographs AbsiMS08 ; Boumal23 ; Sato2021 ; Udri1994 , and articles manopt ; MishS2016 ; SatoA19 ; ShusA2023 ; TiepSon2025 ; WangDPY2024 ; YgerBGR12 , to name a few among many others dedicated to optimization methods on the orthogonal, generalized, and indefinite Stiefel manifolds.
Equipping a Riemannian metric for a differentiable manifold is an essential step in Riemannian optimization. Recall that a Riemannian metric is an inner product defined on each tangent space of the manifold depending smoothly on the contact point. It not only helps quantify objects such as the length of vectors, the measure of angles but also has a great effect on the behavior of numerical methods. Specifically, it is a means of defining the Riemannian gradient of a cost function, a crucial component in most of optimization methods.
There are different ways to equip a manifold with a Riemannian metric depending on how the manifold is viewed. For some left- or right-invariant Riemannian metrics, and pseudo-Riemannian metrics equipped for the orthogonal Stiefel manifold as a quotient manifold of two orthogonal groups, one is referred to (EdelAS98, , Subsect. 2.3), JurdML2020 ; NishA2005 ; cf. BendZ21 for the symplectic Stiefel manifold. Here, we follow TiepSon2025 in which $\mathrm{iSt}_{A,J}(k,n)$ is shown to be an embedded submanifold of $\mathbb{R}^{n\times k}$} of dimension $nk-\frac{1}{2}k(k+1)$. Then, a Riemannian metric is defined via an $n\times n$ spd $X$-smoothly depending matrix $\mathbf{M}_{X}$ in which the inner product of $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}_{A,J}(k,n)$ is determined as
@@
This is not the most general way of defining a metric on a submanifold of $n\times k$ structured matrices but it covers most cases in practice, e.g., the standard Euclidean metric and the canonical metric EdelAS98 . Using this formulation, the Riemannian gradient of smooth cost functions via orthogonally projecting onto the tangent space requires solving a Lyapunov matrix equation (TiepSon2025, , Prop. 3.3), which can be computationally expensive when $k$ is not small. It is surprising that in the case of the orthogonal Stiefel manifold, the Riemannian gradient with respect to the Euclidean and canonical metrics has closed-form expressions; see (EdelAS98, , Sect. 2) and (AbsiMS08, , Subsect. 3.6). Moreover, in the generalized Stiefel manifold $\mathrm{St}_{A}(k,n)$ when $A$ is chosen to be the representing matrix $\mathbf{M}_{X}$, a closed-form Riemannian gradient is available (WangDPY2024, , Sect. 4). Meanwhile, for the symplectic Stiefel manifold, a relatively similar manifold to $\mathrm{St}_{A}(k,n),$ computing the Riemannian gradient needs solving a Lyapunov matrix equation under the Euclidean metric and a general tractable metric and does not when the canonical-like metric is used GSAS21a ; GSAS21 ; GSS2024_2 . On the one hand, these facts raise a hope that one can design a Riemannian metric, most probably the canonical(-like) metric as in EdelAS98 ; GSAS21 , for the indefinite Stiefel manifold such that solving the Lyapunov matrix equation is not compulsory in order to compute the Riemannian gradient. Note, however, that the motivation for using the canonical(-like) metric in EdelAS98 ; GSAS21 is to balance the contribution of the two components of tangent vectors in the metric formulation instead of obviating the necessity of solving the Lyapunov matrix equation as that for the metric proposed in this work. On the other hand, they inspire several questions: (1) How are we on track to avoid solving the Lyapunov matrix equations to compute the Riemannian gradient? (2) How to explain the different situations for orthogonal, generalized, and symplectic Stiefel manifolds mentioned above? And (3) how does an optimization method such as Riemannian gradient descent behave using the geometric quantities under the proposed metric?
@@ 2 Preliminaries
    where $W=JX^{T}AZ,K=(X_{\perp}^{T}A^{-1}X_{\perp})^{-1}X_{\perp}^{T}Z$. It is immediately implied from (4) that $A^{-1}X_{\perp}K=Z-XW$ is independent of $X_{\perp}.$
@@
In (TiepSon2025, , Sect. 3), a so-called tractable metric $\mathrm{g}_{\mathbf{M}_{X}}$ defined via an $X$-smoothly dependent matrix $\mathbf{M}_{X}\in\mathrm{SPD}(n)$ was proposed. Namely, for any pair $Z_{1},Z_{2}\in T_{X}\mathrm{iSt}(k,n),$ $\langle Z_{1},Z_{2}\rangle_{\mathrm{g}_{\mathbf{M}_{X}}}\coloneq\mathrm{tr}\left(Z_{1}^{T}\mathbf{M}_{X}Z_{2}\right)$. With this metric, for any $X\in\mathrm{iSt}(k,n)$, the ambient space $\mathbb{R}^{n\times k}$} is decomposed as the direct sum of the tangent space and the normal space,
@@ 3 A generalized canonical metric on $\mathrm{iSt}(k,n)$
For greater detail, let us convert this matrix equation into a linear equation using vectorization operator and Kronecker product HornJ91 . Namely, for a matrix $M=(m_{ij})\in{\mathbb{R}^{n\times k}}$, $\mathrm{vec}(M)\in\mathbb{R}^{nk}$ is generated by stacking the columns of $M$ to make a column vector, and the Kronecker product of $M$ and $N\in\mathbb{R}^{p\times q}$, denoted by $M\otimes N\in\mathbb{R}^{np\times kq}$, is the matrix consisting of $nk$ block matrices $m_{ij}N$ for $i=1,\ldots,n,j=1,\ldots,k$. In view of (HornJ91, , Lem. 4.3.1), (11) is equivalent to
@@
We say that a square system of linear equations is solving-free if its coefficient is a scalar matrix, i.e., the product of a (nonzero) scalar and the identity matrix. Thus, its solution is merely the product of the reciprocal of the scalar and the right-hand side. Accordingly, a sufficient condition for obviating the need to solve the Lyapunov matrix equation (8) is that $I_{k}\otimes(J\Gamma_{1}J)+(J\Gamma_{1}J)\otimes I_{k}$ has the form $2\rho I_{k^{2}}$, where $\rho$ is a nonzero constant parameter. It follows that $J\Gamma_{1}J=\rho I_{k}$ which, together with the condition $J^{2}=I_{k}$, yields $\Gamma_{1}=\rho I_{k}$. For the symmetric positive-definiteness of $\mathbf{M}_{X}^{-1}$, we set $\rho>0$. One can verify that in this case, the solution to (11) is $U_{X,Y}=\frac{1}{\rho}\mathrm{sym}(X^{T}AY).$ Furthermore, we set $\Gamma_{2}=0$ for the ease of argument. Consequently, $\mathbf{M}_{X}^{-1}$ takes the form
@@
Moreover, taking the formula (3) for the inverse of $E$ into account, we also have
@@
Obviously, $g_{\rho,X_{\perp}}$ is a non-degenerate bilinear form defined on $T_{X}\mathrm{iSt}(k,n)$ and dependent upon both $\rho$ and $X_{\perp}$. To make this a Riemannian metric, it is essential to require that $\mathbf{M}_{X}$ depends only on $X\in\mathrm{iSt}(k,n)$ and this dependence is smooth. Apparently, we can consider $\Gamma_{3}$ as an independent parameter like $\rho$ and smoothly represent $X_{\perp}$ in terms of $X$ (or do this for both $\Gamma_{3}$ and $X_{\perp}$). Nevertheless, we do not know an efficient and explicit formula for this representation, if it exists. Following EdelAS98 ; GSAS21 , we impose a relation on these two components and obtain the desired metric. Namely, we can choose
@@
  • The relations (15) and (17) are only the means to obtain $\mathbf{M}_{X}$. As we will see later, we do not have to construct $\Gamma_{3}$ or $X_{\perp}$.
  • One can see that when $\rho=2$ and $\Gamma_{3}=I_{n-k}$, the formulation (16) reduces to the canonical metric given in (EdelAS98, , Eq. (2.39)). Hence, it is named the generalized canonical metric. Moreover, we use this value for $\rho$ in numerical tests unless stated otherwise.
@@ 3.1 Geometry under the generalized canonical metric
For any $X\in\mathrm{iSt}(k,n),$ the matrix $\mathbf{M}_{X}$ defined by (13) is symmetric positive-definite, so we can extend the inner product $g_{\rho,X_{\perp}}$ on $T_{X}\mathrm{iSt}(k,n)$ to an inner product on $\mathbb{R}^{n\times k}$} which is given by $g_{\rho,X_{\perp}}(Z_{1},Z_{2})=\mathrm{tr}(Z_{1}^{T}\mathbf{M}_{X}Z_{2}),$ for all $Z_{1},Z_{2}\in{\mathbb{R}^{n\times k}}.$ For the moment, neither (15) nor (17) is imposed for the sake of generality. The normal space to $\mathrm{iSt}(k,n)$ at $X$ with respect to $g_{\rho,X_{\perp}}$ is defined by
@@ 3.2 Riemannian gradient
Proof First, recall that the metric $g_{\rho,X_{\perp}}$ in (14) can be extended to $\mathbb{R}^{n\times k}$, and denote by $\mathrm{grad}_{gc}\bar{f}(X)$ the Riemannian gradient of $\bar{f}$ on $\mathbb{R}^{n\times k}$ equipped with that extended metric. In view of the definition of the Riemannian gradient, there holds
@@
for any $Z\in{\mathbb{R}^{n\times k}}.$ Therefore, from the formula (12) for the inverse of $\mathbf{M}_{X}$,
@@
In the case that $\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}X_{\perp}^{T}X_{\perp}\left(X_{\perp}^{T}A^{-1}X_{\perp}\right)^{-1}=\Gamma_{3}$, taking the formula (5) resulting from the identity $I_{n}=EE^{-1}$ into account, it follows that
@@
If matrix $A$ is dense and has no additional special structure, the computational cost for $A^{-1}(\cdot)$ in (23) is generally comparable to $n^{3}$. Therefore, in this case, using (24) is more advantageous. One can see that the computation of (24) requires $2n^{2}k+10nk^{2}-nk-3k^{2}$ flops. Obviously, when $k$ is of the same scale as $n$, i.e., $k=\mathcal{O}(n)$, this cost is cheaper than that caused by (9). This will be seen in the numerical examples presented in Section 6.
@@
In spite of the difference, similar arguments also work for the geometry of the symplectic Stiefel manifold under the canonical-like metric GSAS21 as a special case of the tractable metric GSS2024_2 .